(5 Points) Use the unit circle values to find the value of the trigonometric function. Show all work! csc (****)
Answer #1

Solution:

A unit circle is a circle with radius r = 1 unit

\small csc\left (- \frac{37\pi }{4} \right )

\small \Rightarrow csc\left (- \frac{36\pi }{4}-\frac{\pi }{4} \right )

\small \Rightarrow \frac{1}{sin\left ( -9\pi -\frac{\pi }{4} \right )}

So, \small \theta take 4 complete revolution and 1 half revolution in anti-clockwise direction, then it moves pi/4 in same direction. \small \theta stops at \small 3\pi /4 position in unit circle.

since  \small sin\frac{3\pi }{4} = \frac{\sqrt{3}}{2}

Hence,

\small \Rightarrow csc\left ( -\frac{37\pi }{4} \right )= \frac{1}{sin\left ( -9\pi -\frac{\pi }{4} \right )} = \frac{2}{\sqrt{3}}

\small \Rightarrow csc\left ( -\frac{37\pi }{4} \right ) = \frac{2}{\sqrt{3}}

XUnit Circle 4号) (0.1) 90°、 (3) 6 09 (59) cose,sino) ( 学) 120- | 1353, 3 150/50 \ 1- 3 (-1,0), soln - - 0 s (10) 2lol /330

PS: hit the like button